5. A soapbox derby race car starts at rest and moves 300 m down a track in 22.4 s. Calculate its acceleration (assumed constant) and its speed at the end of the track.
1
Expert's answer
2020-12-10T11:09:47-0500

The distance travelled in time "t" under the constant acceleration (and with zero initial speed) is:


"d = \dfrac{at^2}{2}"

where "a" is the acceleration and "t = 22.4s" is the time. If the distance is "d = 300m", then expressing "a", obtain:


"a = \dfrac{2d}{t^2} = \dfrac{2\cdot200m }{(22.4s)^2} \approx 0.8m/s^2"

The speed at the end of the track is:


"v = \dfrac{2d}{t^2}t = \dfrac{2d}{t}\\
v = \dfrac{2\cdot 200m}{22.4s} \approx 17.9m/s"

Answer. 0.8 m/s^2, 17.9 m/s.